a soil analysis of a lawn determines that 50 pounds of fertilizeer containing 20 % nitrogen is needed. How can this mixture be made from two different fertilizers, one containing 15% nitrogen and the other containing 24% nitrogen?
Answer by josmiceli(19441)

x = 15% fertilizer
y = 24% fertilizer
x + y = 50
.15x + .24y = .20*50
.15x + .24(50 - x) = 10
.15x + 12 -.24x = 10
- .09x = -2
x = 22 2/9 (pounds of 15% fertilizer)
y = 50 - 22 2/9
y = 27 7/9 (pounds of 24% fertilizer)